# reference.bib

### v1

@@ Abstract
The rise of the Internet of Things and edge computing has shifted computing resources closer to end-users, benefiting numerous delay-sensitive, computation-intensive applications. To speed up computation, distributed computing is a promising technique that allows parallel execution of tasks across multiple compute nodes. However, current research predominantly revolves around the master-worker paradigm, limiting resource sharing within one-hop neighborhoods. This limitation can render distributed computing ineffective in scenarios with limited nearby resources or constrained/dynamic connectivity. In this paper, we address this limitation by introducing a new distributed computing framework that extends resource sharing beyond one-hop neighborhoods through exploring layered network structures and multi-hop routing. Our framework involves transforming the network graph into a sink tree and formulating a joint optimization problem based on the layered tree structure for task allocation and scheduling. To solve this problem, we propose two exact methods that find optimal solutions and three heuristic strategies to improve efficiency and scalability. The performances of these methods are analyzed and evaluated through theoretical analyses and comprehensive simulation studies. The results demonstrate their promising performances over the traditional distributed computing and computation offloading strategies.
@@ I Introduction
In this paper, we overcome these challenges by exploring resources at distant servers located multiple hops away. While a similar idea has been explored in the mobile edge computing (MEC) domain, where computation offloading is proposed to address users’ computing demands, most existing studies focus on offloading to a single MEC server [kuang2019partial, tang2020partial, 9435782, you2016energy, barbarossa2013joint, li2018deep]. A few recent works [zhang2022result, xie2022dynamic, zhang2022joint, funai2019computational] have considered offloading tasks to multiple servers, but they overlook the task scheduling issue addressed in this work and offer only heuristic solutions. To the best of our knowledge, this is the first systematic investigation into computation offloading to multiple servers that can be multiple hops away, and we propose efficient exact solutions for optimal task allocation and scheduling.
@@
  • A new multi-hop multi-layer distributed computing framework: The proposed framework explores layered network structures and multi-hop routing to fully utilize the capacity of the entire edge computing network for enhanced system performance. It transforms the network graph into a sink tree, based on which a mixed integer programming (MIP) problem is formulated to jointly optimize task allocation and scheduling. Notably, this framework can be applied to any networked computing systems such as cloud computing, MEC, and networked airborne computing systems.
  • Two exact methods that find optimal solutions: To solve the optimization problem, we propose two exact methods, including a centralized method and a parallel enhancement of it. We also present an offline-online computation scheme that allows both methods to execute in real-time and handle dynamic and mobile networks.
@@
The rest of the paper is organized as follows. Sec. II discusses related works. Sec. III describes the system model and the problem to be solved. Sec. IV introduces the proposed multi-hop multi-layer distributed computing framework. The two exact methods and the three heuristic methods are introduced in Sec. V and Sec. VI, respectively. Sec. VII presents the results of simulation studies. Finally, Sec. VIII concludes the paper and discusses the future works.
@@ II-B Computation Offloading
In the computing offloading domain, most existing studies consider a single-hop single-server offloading paradigm, where tasks are offloaded from users to a single edge server within their communication range [kuang2019partial, tang2020partial, li2018deep, 9435782, barbarossa2013joint, you2016energy]. The tasks can be offloaded as a whole or partially, known as binary offloading and partial offloading, respectively. Under this paradigm, many algorithms have been designed to make the optimal offloading decisions. For instance, studies in [kuang2019partial, tang2020partial] examine the scenario where tasks are offloaded from a single user to a single nearby server. There have also been studies that explore tasks from multiple users, optimizing not only the offloading decisions (whether to offload a task or determining the offloading ratio) but also the allocation of resources to each user. For instance, [li2018deep] considers the allocation of computing resources, while [9435782, barbarossa2013joint] addresses the allocation of both computing and transmission power resources. The allocation of communication resources, including time slots under the Time Division Multiple Access protocol and sub-channels under the Orthogonal Frequency-Division Multiple Access protocol, is considered in [you2016energy]. These problems are typically solved using numerical approaches [you2016energy, barbarossa2013joint]. Reinforcement learning has also emerged as a promising tool for computation offloading [9435782, li2018deep].
@@
In practical scenarios, it is possible that there are no (powerful) edge servers nearby for the end users. To address this limitation, researchers have started to explore multi-hop offloading [mach2017mobile], enabling the offloading of tasks from users to remote servers multiple hops away. Along this direction, existing studies mostly consider offloading tasks to a single server, such as the studies in [hong2019multi, liu2022mobility, wang2020energy]. There are also several works that investigate partitioning tasks into multiple parts and offloading these parts to multiple servers, which are most relevant to our study [zhang2022result, zhang2022joint, funai2019computational, xie2022dynamic]. For instance, [zhang2022result] investigates the joint routing and multi-part offloading for both data and result. It employs a flow model to capture data/result traffic and introduces a distributed algorithm that finds optimal solutions in polynomial time. [xie2022dynamic] formulates the multi-hop offloading problem as a potential game. By dividing tasks into subtasks of equal size, each device independently decides the number of subtasks to forward or compute based on its economic utility. The study in [funai2019computational] addresses the distribution of a set of tasks, partitioned from a complex application, to multiple cooperative servers that may be multiple hops away. This problem is formulated as a task assignment problem and solved by an iterative algorithm. Another relevant work is presented in [zhang2022joint], which considers a joint user association, channel allocation, and task offloading problem. It solves this problem by combining the genetic algorithm and deep deterministic policy gradient algorithm.
@@
Distinct from previous research, we delve into the essential benefits of multi-hop routing and multi-part offloading while investigating how network properties like topology and server resources affect system performance. We also address the task scheduling problem that arises when transmissions of subtasks share channels or relays, which has been overlooked by existing works. Moreover, we propose both exact and heuristic methods to solve the problem, and introduce an offline-online computation scheme to enable real-time implementation and make them adaptable to dynamic and mobile networks.
@@ III-A System Model
Consider a network formed by $N+1$ edge servers, each with its own unique set of computing and communication capabilities. The servers can share resources with their neighbors through cables (in wired networks) or when they are within communication range (in wireless networks). The entire system is supervised and managed by a control center (e.g., a software defined networking controller [kreutz2014software]) to ensure that all tasks are completed efficiently and effectively.
Suppose one of the servers, referred to as master, needs to execute a computation-intensive task that is arbitrarily decomposable, which could be generated by the server itself or requested by a user nearby. To complete the task in a timely and energy-efficient manner, the master decomposes the task into subtasks and distributes them to the other servers, referred to as workers (see Fig. 1 for an illustration). The master (highlighted with red) can transmit subtasks simultaneously to their neighboring servers. In the context of wireless communication, we assume the use of the Orthogonal Frequency Division Multiple Access technology [zhang2022joint]. However, for workers farther away, multi-hop routing is required, which means that each server in the network can act as a worker, a relay, or both. When a subtask arrives at a relay, it is added to a queue and processed in a first-in-first-out order. A worker will not start executing the assigned subtask until it receives the complete subtask package. When a server acts as both a worker and a relay, it can perform the relay process and execute the assigned task simultaneously.
@@ III-A3 Communication Model
Let $R_{i,j}$ be the data transmission rate from server $i$ to server $j$, which can be obtained by the following equation according to Shannon’s Theory [zhang2023exploring]:
$R_{i,j}$ $=B_{i,j}log_{2}(1+\frac{s_{i,j}}{n_{i,j}})$ (2)
where $B_{i,j}$ denotes the bandwidth of the communication link $(i,j)$. $s_{i,j}$ and $n_{i,j}$ represent the signal power and noise power, respectively.
@@ III-B Problem Description and Analysis
IV Multi-Hop Multi-Layer Distributed Computing Framework
In this section, we present a multi-hop multi-layer distributed computing framework to solve the problem described in the previous section. Motivated by the fact that a layered tree structure emerges when the master distributes tasks to other servers in the network, this framework first transforms the network graph into a sink tree and exploits this layered tree structure to find optimal task allocation and scheduling solutions.
@@ IV-A Transforming Graph into a Sink Tree
To facilitate subsequent analysis, we re-label the nodes in the tree $\mathcal{T}$ level-by-level from the root downward, and from left to right within each level (see Fig. 2). Consequently, nodes in lower levels have larger indices. Let $\mathcal{I}_{l}$ denote the set of indices of nodes in level $l\in\{0,1,\ldots,H\}$, where $H$ is the height of the tree. Then, $\cup_{l=0}^{H}\mathcal{I}_{l}=\mathcal{N}$. Notably, the master (root) can transmit subtasks to its one-hop neighbors, i.e., nodes in Level $1$, simultaneously using orthogonal channels. However, if any one-hop neighbor has children, the subtasks assigned to them, including the one-hop neighbor, have to be transmitted one by one. This is because they share the same channel between the master and the one-hop neighbor, and the data arriving at the one-hop neighbor is processed in a first-in-first-out manner. Therefore, the order in which these subtasks should be sent matters. Based on these analyses, we next formulate a joint task allocation and scheduling problem as a mixed integer programming (MIP) model.
@@ VI-A2 Level Pruning (LP)
The key idea of LP is to trim nodes that are excessively distant from the master node, whose computing resources are too costly to use considering the significant communication costs. Specifically, this strategy evaluates the top $\xi+1$ levels of the original network tree, removing levels from $\xi$ to $H$. The resulting tree, denoted as $\mathcal{T}^{\xi}$, satisfies $\mathcal{T}^{\xi}=\mathcal{T}\setminus\cup_{l=\xi}^{H}\mathcal{I}_{l}$.
@@ VI-B Genetic Algorithm
The worker selection methods allow us to reduce workers but may prune nodes that could significantly improve system performance. Here, we introduce a genetic algorithm (GA) [mirjalili2019genetic] that allows us to evaluate large networks. The complete procedure is outlined in Algorithm 3. It involves two phases: initialization and training. In the initialization phase (Lines 1-5), a population set $\boldsymbol{O}=\{\boldsymbol{o}_{k}\}$ is first randomly generated, which consists of $P$ offloading orders (chromosomes). The corresponding optimal task partition $\boldsymbol{y}^{*}(\boldsymbol{o}_{k})$ and cost $z^{*}(\boldsymbol{o}_{k})$ are then computed by solving (12), where $z^{*}(\boldsymbol{o}_{k})$ is the fitness of the chromosome $\boldsymbol{o}_{k}$. Following the initialization, the training phase (Lines 6-17) starts with Elitism, which picks the top $\alpha\%$ of the fittest members from the current population and propagates them to the next generation. After that, an iterative procedure is performed to create offspring. In each iteration, two offloading orders are randomly picked from the current population $\boldsymbol{O}$ according to the probability distribution $\boldsymbol{a}$ computed in Line 8. Ordered crossover [deep2011new] is then applied to create offspring. Subsequently, with a low probability $\beta$, mutation is performed to introduce diversity into the new population by shuffling individual offloading orders. The algorithm terminates upon meeting the stopping condition at which point it outputs the best solution found. In our simulations, we set the stopping condition for GA to be reaching a maximum number of generations, denoted as $G$.
Algorithm 3 GA($\mathcal{T}$, $Y$)
1:  $\boldsymbol{O}\leftarrow\varnothing$;
2:  for $k\leftarrow 1$ to $P$ do
3:     $\boldsymbol{o}_{k}\leftarrow$ a randomly generated offloading order;
4:     Compute $\boldsymbol{y}^{*}(\boldsymbol{o}_{k}),z^{*}(\boldsymbol{o}_{k})$ using (12);
5:  end for
6:  for $n\leftarrow 1$ to $G$ do
7:     $\boldsymbol{O}^{\prime}\leftarrow$ top $\alpha$of offloading orders in set $\boldsymbol{O}$ with the smallest costs;
8:     $\boldsymbol{a}\leftarrow\left\{\left.\frac{1/z^{*}(\boldsymbol{o}_{k})}{\sum_{j=1}^{P}1/z^{*}(\boldsymbol{o}_{j})}\right|k\in\{1,2,\ldots,P\}\right\}$;
9:     while $|\boldsymbol{O}^{\prime}|<P$ do
10:        $\boldsymbol{o}_{i},\boldsymbol{o}_{j}\leftarrow$ two offloading orders randomly selected from $\boldsymbol{O}$ according to the probability distribution $\boldsymbol{a}$;
11:        $\boldsymbol{o}^{\prime}_{i},\boldsymbol{o}^{\prime}_{j}\leftarrow$ crossover($\boldsymbol{o}_{i},\boldsymbol{o}_{j}$);
12:        With a probability of $\beta$, apply mutation to $\boldsymbol{o}^{\prime}_{i}$, $\boldsymbol{o}^{\prime}_{j}$;
13:        $\boldsymbol{O}^{\prime}\leftarrow\boldsymbol{O}^{\prime}\cup\{\boldsymbol{o}^{\prime}_{i},\boldsymbol{o}^{\prime}_{j}\}$;
14:        Compute $\boldsymbol{y}^{*}(\boldsymbol{o}_{i}),z^{*}(\boldsymbol{o}_{i})$ and $\boldsymbol{y}^{*}(\boldsymbol{o}_{j}),z^{*}(\boldsymbol{o}_{j})$ using (12);
15:     end while
16:     $\boldsymbol{O}\leftarrow\boldsymbol{O}^{\prime}$;
17:  end for
18:  $\boldsymbol{o}^{\star}\leftarrow\operatorname*{arg\,min}_{\boldsymbol{o}_{k}\in\boldsymbol{O}}z^{*}(\boldsymbol{o}_{k})$, $y^{\star}\leftarrow y^{*}(\boldsymbol{o}^{\star})$, $z^{\star}\leftarrow z^{*}(\boldsymbol{o}^{\star})$;
19:  return  $\boldsymbol{o}^{\star}$, $\boldsymbol{y}^{\star}$, $z^{\star}$;

### v2

@@ Abstract
The rise of the Internet of Things and edge computing has shifted computing resources closer to end-users, benefiting numerous delay-sensitive, computation-intensive applications. To speed up computation, distributed computing is a promising technique that allows parallel execution of tasks across multiple compute nodes. However, current research predominantly revolves around the master-worker paradigm, limiting resource sharing within one-hop neighborhoods. This limitation can render distributed computing ineffective in scenarios with limited nearby resources or constrained/dynamic connectivity. In this paper, we address this limitation by introducing a new distributed computing framework that extends resource sharing beyond one-hop neighborhoods through exploring layered network structures. Our framework involves transforming the network graph into a sink tree and formulating a joint optimization problem based on the layered tree structure for task allocation and scheduling. To solve this problem, we propose two exact methods that find optimal solutions and three heuristic strategies to improve efficiency and scalability. The performances of these methods are analyzed and evaluated through theoretical analyses and comprehensive simulation studies. The results demonstrate their promising performances over the traditional distributed computing and computation offloading strategies.
@@ I Introduction
In this paper, we overcome these challenges by exploring resources at distant servers located multiple hops away. While a similar idea has been explored in the mobile edge computing (MEC) [kuang2019partial, tang2020partial, liu2023rfid, 9435782, you2016energy, barbarossa2013joint, li2018deep] and Internet of Vehicle [chen2022multihop, liu2022mobility, zhao2024asynchronous, deng2020multi] domains, where computation offloading is proposed to address users’ and vehicles’ computing demands, most existing studies focus on offloading tasks to a single MEC server located one or multiple hops away, with intermediate servers serving solely as relays. A few recent works [zhang2022result, xie2022dynamic, zhang2022joint, funai2019computational] have considered offloading tasks to multiple servers, but they overlook the task scheduling issue addressed in this work and offer only heuristic solutions. To the best of our knowledge, this is the first systematic investigation into computation offloading to multiple servers across multiple hops and into the benefits of layered structures for improving distributed computing performance. Additionally, we present efficient exact solutions for optimal task allocation and scheduling, which generalize existing solutions and are applicable to a wide range of networked computing networks.
@@
  • A new multi-layered distributed computing framework: The proposed framework explores layered network structures to fully utilize the capacity of the entire edge computing network for enhanced system performance. It transforms the network graph into a sink tree, based on which a mixed integer programming (MIP) problem is formulated to jointly optimize task allocation and scheduling. Notably, this framework can be applied to any networked computing systems such as cloud computing, MEC, and networked airborne computing systems.
  • Two exact methods that find optimal solutions: To solve the optimization problem, we propose two exact methods, including a centralized method and a parallel enhancement of it. We also present an offline-online computation scheme that allows both methods to execute in real-time and handle dynamic and mobile networks. How these methods generalize existing solutions is also discussed.
@@
It should be noted that the multi-layered distributed computing framework was initially introduced in a short conference version [10622383], which presented a different heuristic method for solving the MIP problem. This paper provides a more comprehensive and systematic investigation, featuring a set of new and rigorously analyzed methods.
The rest of the paper is organized as follows. Sec. II discusses related works. Sec. III describes the system model and the problem to be solved. Sec. IV introduces the proposed multi-layered distributed computing framework. The two exact methods and the three heuristic methods are introduced in Sec. V and Sec. VI, respectively. Sec. VII presents the results of simulation studies. Finally, Sec. VIII concludes the paper and discusses the future works.
@@ II-B Computation Offloading
In the computing offloading domain, most existing studies consider a single-hop single-server offloading paradigm, where tasks are offloaded from users to a single edge server within their communication range [kuang2019partial, tang2020partial, liu2023rfid, li2018deep, 9435782, barbarossa2013joint, you2016energy]. The tasks can be offloaded as a whole or partially, known as binary offloading and partial offloading, respectively. Under this paradigm, many algorithms have been designed to make the optimal offloading decisions. For instance, studies in [kuang2019partial, tang2020partial] examine the scenario where tasks are offloaded from a single user to a single nearby server. [liu2023rfid] extends this analysis by taking server mobility and task dependency into account. There have also been studies that explore tasks from multiple users, optimizing not only the offloading decisions (whether to offload a task or determining the offloading ratio) but also the allocation of resources to each user. For instance, [li2018deep] considers the allocation of computing resources, while [9435782, barbarossa2013joint] addresses the allocation of both computing and transmission power resources. The allocation of communication resources, including time slots under the Time Division Multiple Access protocol and sub-channels under the Orthogonal Frequency-Division Multiple Access (OFDMA) protocol, is considered in [you2016energy]. These problems are typically solved using numerical approaches [you2016energy, barbarossa2013joint]. Reinforcement learning has also emerged as a promising tool for computation offloading [9435782, li2018deep].
In practical scenarios, it is possible that there are no (powerful) edge servers nearby for the end users. To address this limitation, researchers have started to explore multi-hop offloading, enabling the offloading of tasks from users to remote servers multiple hops away. Along this direction, existing studies mostly consider offloading tasks to a single server, as seen in [hong2019multi, wang2020energy, chen2022multihop, liu2022mobility, zhao2024asynchronous, deng2020multi]. Additionally, the three-tier network topology comprising end users, edge servers and cloud servers [qi2024bridge, liu2023joint, sun2023vehicular, almuseelem2023energy] has garnered considerable interest. In their approach, tasks are offloaded either to a nearby edge server one hop away or to a cloud server two hops away, with edge servers acting as relays. In contrast, we consider all servers reachable by users and aim to facilitate collaborative computing among them.
@@
There are also several works that investigate partitioning tasks into multiple parts and offloading these parts to multiple servers, which are most relevant to our study [zhang2022result, zhang2022joint, funai2019computational, xie2022dynamic]. For instance, [zhang2022result] investigates the joint routing and multi-part offloading for both data and result. It employs a flow model to capture data/result traffic and introduces a distributed algorithm that finds optimal solutions in polynomial time. [xie2022dynamic] formulates the multi-hop offloading problem as a potential game. By dividing tasks into subtasks of equal size, each device independently decides the number of subtasks to forward or compute based on its economic utility. The study in [funai2019computational] addresses the distribution of a set of tasks, partitioned from a complex application, to multiple cooperative servers that may be multiple hops away. This problem is formulated as a task assignment problem and solved by an iterative algorithm. Another relevant work is presented in [zhang2022joint], which considers a joint user association, channel allocation, and task offloading problem. It solves this problem by combining the genetic algorithm and deep deterministic policy gradient algorithm.
@@
Distinct from previous research, we delve into the essential benefits of layered network structures while investigating how network properties like topology and server resources affect system performance. We also address the task scheduling problem that arises when transmissions of subtasks share channels or relays, which has been overlooked by existing works. Moreover, we propose both exact and heuristic methods to solve the problem, and introduce an offline-online computation scheme to enable real-time implementation and make them adaptable to dynamic and mobile networks.
@@ III-A System Model
Consider a network formed by $N+1$ edge servers, each with its own unique set of computing and communication capabilities. The servers can share resources with their neighbors either through cables in wired networks or wirelessly when they are within communication range. Additionally, a server can communicate with its one-hop neighbors simultaneously using techniques like Orthogonal Frequency Division Multiplexing [hwang2008ofdm]. For simplicity, interference among the servers is not considered in this study. The entire system is supervised and managed by a control center (e.g., a software defined networking controller [kreutz2014software]) to ensure that all tasks are completed efficiently and effectively.
Suppose one of the servers, referred to as master, needs to execute a computation-intensive task that is arbitrarily decomposable, which could be generated by the server itself or requested by a user nearby. To complete the task in a timely and energy-efficient manner, the master decomposes the task into subtasks and distributes them to the other servers, referred to as workers (see Fig. 1 for an illustration). The master (highlighted with red) can transmit subtasks simultaneously to their neighboring servers. However, for workers farther away, multi-hop routing is required, which means that each server in the network can act as a worker, a relay, or both. When a subtask arrives at a relay, it is added to a queue and processed in a first-in-first-out order. A worker will not start executing the assigned subtask until it receives the complete subtask package. When a server acts as both a worker and a relay, it can perform the relay process and execute the assigned task simultaneously.
@@ III-A3 Communication Model
Let $R_{i,j}$ denote the data transmission rate from server $i$ to server $j$, which we assume to be known for the sake of simplifying analysis. This rate can be approximated using the Shannon’s Theory [zhang2023exploring] as $R_{i,j}=B_{i,j}log_{2}(1+\frac{s_{i,j}}{n_{i,j}})$, where $B_{i,j}$ is the channel bandwidth, and $s_{i,j}$ and $n_{i,j}$ represent the signal power and noise power, respectively.
@@ III-B Problem Description and Analysis
IV Multi-Layered Distributed Computing Framework
In this section, we present a multi-layered distributed computing framework to solve the problem described in the previous section. Motivated by the fact that a layered tree structure emerges when the master distributes tasks to other servers in the network, this framework first transforms the network graph into a sink tree and exploits this layered tree structure to find optimal task allocation and scheduling solutions.
@@ IV-A Transforming Graph into a Sink Tree
To facilitate subsequent analysis, we re-label the nodes in the tree $\mathcal{T}$ level-by-level from the root downward, and from left to right within each level (see Fig. 2). Consequently, nodes in lower levels have larger indices. Let $\mathcal{I}_{l}$ denote the set of indices of nodes in level $l\in\{0,1,\ldots,H\}$, where $H$ is the height of the tree. Then, $\cup_{l=0}^{H}\mathcal{I}_{l}=\mathcal{N}$. Notably, the master (root) can transmit subtasks to its one-hop neighbors, i.e., nodes in Level $1$, simultaneously. However, if any one-hop neighbor has children, the subtasks assigned to them, including the one-hop neighbor, have to be transmitted one by one. This is because they share the same channel between the master and the one-hop neighbor, and the data arriving at the one-hop neighbor is processed in a first-in-first-out manner. Therefore, the order in which these subtasks should be sent matters. Based on these analyses, we next formulate a joint task allocation and scheduling problem as a mixed integer programming (MIP) model.
@@ VI-A2 Level Pruning (LP)
The key idea of LP is to trim nodes that are excessively distant from the master node, whose computing resources are too costly to use considering the significant communication costs. Specifically, this strategy evaluates the top $\xi$ levels of the original network tree, removing levels from $\xi+1$ to $H$. The resulting tree, denoted as $\mathcal{T}^{\xi}$, satisfies $\mathcal{T}^{\xi}=\mathcal{T}\setminus\cup_{l=\xi}^{H}\mathcal{I}_{l}$.
@@ VI-B Genetic Algorithm
The worker selection methods allow us to reduce workers but may prune nodes that could significantly improve system performance. Here, we introduce a genetic algorithm (GA) [mirjalili2019genetic] that allows us to evaluate large networks. It involves two phases: initialization and training. In the initialization phase, a population set $\boldsymbol{O}=\{\boldsymbol{o}_{k}\}$ is first randomly generated, which consists of $P$ offloading orders (chromosomes). The corresponding optimal task partition $\boldsymbol{y}^{*}(\boldsymbol{o}_{k})$ and cost $z^{*}(\boldsymbol{o}_{k})$ are then computed by solving (11), where $z^{*}(\boldsymbol{o}_{k})$ is the fitness of the chromosome $\boldsymbol{o}_{k}$. Following the initialization, the training phase starts with Elitism, which picks the top $\alpha\%$ of the fittest members from the current population and propagates them to the next generation. After that, an iterative procedure is performed to create offspring. In each iteration, two offloading orders are randomly picked from the current population $\boldsymbol{O}$ according to the probability distribution $\left\{\left.\frac{1/z^{*}(\boldsymbol{o}_{k})}{\sum_{j=1}^{P}1/z^{*}(\boldsymbol{o}_{j})}\right|k\in\{1,2,\ldots,P\}\right\}$. Ordered crossover [deep2011new] is then applied to create offspring. Subsequently, with a low probability $\beta$, mutation is performed to introduce diversity into the new population by shuffling individual offloading orders. The algorithm terminates upon meeting the stopping condition at which point it outputs the best solution found. In our simulations, we set the stopping condition for GA to be reaching a maximum number of generations, denoted as $G$.